OpenStudy (annie96):

Help Do I have this answers right. Each small Kite cost $4, while each large Kite cost $7. If the total money collected selling Kites is $150, and 30 Kites are sold, how many of each size Kite are sold? A) 8 small and 22 large B) 10 small and 20 large (Answer) C) 20 small and 10 large D) 27 small and 6 large Which system has no solutions? A) x<5 x>1 B) x>5 x<1 C) x<5 x<1 D) x>5 x>1

OpenStudy (anonymous):

How did you get your answers?

OpenStudy (anonymous):

Unfortunately, they both seem to be wrong.

OpenStudy (anonymous):

For linear combinations you must solve 1 equation for 1 of the variables( for example x or y) and then use the solution of your variable you solved for and plug it back it into the second equation and solve for the variable left to get an answer for 1 and from there you can find the other.

OpenStudy (anonymous):

first one should be, 0,4)

OpenStudy (anonymous):

3x + y = 4 I am going to rearrange to y =ax + b form y = -3x + 4 sub y = -3x + 4 into second equation 2x + y = 4 2x + (-3x + 4) = 4 2x -3x + 4 = 4 -x = 0 x = 0 sub x = 0 back into either equations to find y 2x + y = 4 2(0) + y = 4 0 + y = 4 y = 4

OpenStudy (annie96):

ok

OpenStudy (anonymous):

@Johnbc wanna do 2 ? :D